# Exploring the Limitations of Graph Reasoning in Large Language Models

### v1

@@ Abstract
Pretrained Large Language Models have demonstrated various types of reasoning capabilities through language-based prompts alone. However, in this paper, we test the depth of graph reasoning for 5 different LLMs (GPT-4, GPT-3.5, Claude-2, Llama-2 and Palm-2) through the problems of graph reasoning. In particular, we design 10 distinct problems of graph traversal, each representing increasing levels of complexity. Further, we analyse the performance of models across various settings such as varying sizes of graphs as well as different forms of k-shot prompting. We highlight various limitations, biases and properties of LLMs through this benchmarking process, such as an inverse relation to the average degrees of freedom of traversal per node in graphs, the overall negative impact of k-shot prompting on graph reasoning tasks, and a positive response bias which prevents LLMs from identifying the absence of a valid solution. Finally, we propose a new prompting technique specially designed for graph traversal tasks, known as PathCompare, which shows a notable increase in the performance of LLMs in comparison to standard prompting and CoT.
Index Terms:
Large Language Models, LLMs, reasoning, graph, benchmark
 [^1] All code for API-based prompting, as well as results, can be found on our github repository.
I Introduction
Large language models (LLMs) have made significant advances in recent years, demonstrating impressive natural language processing capabilities. These models can generate human-like text and engage in layered contextual conversation. This has led to excitement about the potential for LLMs to exhibit intelligent behavior and assist humans in a variety of tasks involving language and reasoning. However, there is debate around whether the latest iterations of LLMs actually “understand” language or display true reasoning abilities. While they can certainly generate coherent text, it is unclear whether LLMs construct valid underlying relation graphs that carry true logical reasoning.
Many works have explored the various reasoning capabilities of LLMs [1], including arithmetic [2], logical [3], semantic [4], and symbolic [5]. Various techniques,such as chain-of-thought and self-consistency[6], have been proposed to improve the reasoning abilities of LLMs by enhancing the implicit knowledge in the output.
In this paper, we are particularly interested in studying the graph reasoning abilities. Graphs represent a particular aspect of reasoning that cannot be easily captured by text-based or arithmetic reasoning. Graph reasoning can cover non-linear connectivity between multi-dimensional elements in a condensed format. Traversal over graphs requires simultaneous tracking of the state of multiple nodes, which may be somewhat challenging for LLMs. Through a structured formulation of graph problems, LLMs can be systematically tested on their analytical reasoning capabilities, leading to some interesting insights. Particularly, graph reasoning is a proxy for understanding how LLMs establish relations between different entities.
Some works have modeled LLMs to perform specific graph reasoning tasks. In [7], a graph transformer is used to encode a graph, the encodings from which are used to fine-tune an LLM. However, this defeats the purpose of understanding emerging reasoning abilities as a result of pretraining of large corpora of data. Fine-tuning results represent the convergence abilities of the model training process, rather than the effect of large scale pretraining on the reasoning abilities of LLMs.
@@ I Introduction
A systematic graph-reasoning benchmarking of pretrained LLMs was performed in [8], where various interesting properties were revealed. This included preliminary reasoning capabilities in LLMs in basic graph traversal problems, and the ineffectual nature of advanced prompting techniques in the context of logical deduction. In the aforementioned work, an improvement in LLM reasoning is observed as a result of custom prompting, majorly along the lines of directing the LLM to elaborate on graph connections in order to facilitate node traversal. While this is a comprehensive study on its own, we build on top of this work to study the problem through a wider and more granular approach. Specifically, we carry out this evaluation through a series of increasingly complex graph problems and settings. Particularly, we construct 10 distinct graph problems requiring multi-hop reasoning and tracking. This includes tree-based graph traversals, grid-based graph-traversals as well as certain special classes of problems. We evaluate five different LLMs (GPT-3.5, GPT-4, Claude-2, Llama-2 and Palm-2). While the OpenAI family of models is thoroughly investigated in various papers, this paper studies the nature of various other contenders. For example, in this paper, we demonstrate that Anthropic’s Claude-2 is a good logical reasoning model, only sub-standard to GPT-4 among the various models publically available for access. We also choose to adopt certain design choices that highlight interesting limitations and biases in the training of various models. This includes methods like jumbling the sequence of nodes, as well as testing if the models can identify the absence of a valid solution, or in other words, testing if the models are biased to responding with a non-negative response.
@@
    We benchmark five different LLM models on graph reasoning tasks. The graph reasoning tasks are carefully designed in the form of 10 traversal problems in increasing orders of complexity.
@@
    Further, to examine models incrementally in a systematic manner, we compare results across three k-shot settings as well as three different graph order complexities. For each setting, 10 examples are aggregated and averaged for statistical consistency. Through this, we separate models with notable reasoning abilities from models which are in essence stochastic token generators.
@@
    We compare various properties of LLMs through special metrics (i.e. partial accuracy, versus the standard binary accuracy), which gives more granular insights into the reasoning abilities of LLMs.
  4. 4.
    We finally propose a new prompting technique which improves graph traversal performance across different LLMs. We demonstrate that this prompting technique outperforms standard prompting as well as CoT prompting in the majority (i.e. $>$50%) of the designed tasks).
@@
II Experimental setting
@@
II-A Defining graph levels and complexity
@@
these graphs are defined as connected undirected graphs, where any two nodes $u,v$ are connected by atmost one edge $e\epsilon\{0,R^{+}\}$, (where $R^{+}$ represents positive real number values). In the above representation $e=0$ represents no connection, $e=1$ represents a simple unweighted connection between nodes $u,v$, and all other positive values $e~{}\epsilon~{}(R^{+}|e\neq 1$) represents a weighted connection between nodes $u,v$. By definitions, no cycles exist in a tree-based graph. However, for our purpose, we define tree-based graphs as tree-like heirarchical structures with possible exceptions. The emphasis is on navigating through these arbitrary node connections to find the shortest or least costly route. Based on this, we formulate the following problems
@@
II-B Graph generation, prompting and evaluation
@@
II-C Selection of LLMs
In order to include a broad range of models, we select five models from four families of models (in terms of architecture and training data/method). This includes – OpenAI’s GPT-3.5 and GPT-4 [10], Anthropic’s Claude-2 [11], Google’s Palm-2 [12], and Meta’s Llama-2 [13] with 13B parameters. Since LLama-2 is a raw foundational model, we use an instruction-fine tuned version (Hermes) released by Nous Research.
III General Discussion
Using the prompting and evaluation technique defined above, we evaluate the various models over the various problems. We evaluate the results of graph reasoning for problem 1, problem 2, and problem 3, by calculating averaged and normalized binary accuracies across 10 examples for each setting, and go from 0 to 1 in increments of 0.1. A model response is only marked correct if it predicted the shortest/least-cost traversal path correctly and completely. Along with binary accuracies, partial accuracies are also calculated. While we use binary accuracy to filter out models with weak reasoning abilities, we use partial accuracy for more granular comparison.
@@ III General Discussion
TABLE I: Evaluation of models on problem category 1: tree-based graph traversals. The values are represented as A/B, where A depicts the average normalized binary accuracy (ranging from 0 to 1) for 10 examples per setting, and B depicts the average normalized value for partial credit for 10 examples per setting. * = a larger context variant of the corresponding LLM was used. $\bigtriangleup$ = the prompt was too large to fit the context window of the model, hence the setting was omitted. ** = the model did not return any solution and refused to answer the question.
@@
TABLE II: Evaluation of models on problem category 2: grid-based graph traversals. The values are represented as A/B, where A depicts the average normalized binary accuracy (ranging from 0 to 1) for 10 examples per setting, and B depicts the average normalized value for partial credit for 10 examples per setting. The notations are similar to Table I.
@@
TABLE III: Evaluation of models on problem category 3: special problems. The values depict the average normalized value (ranging from 0 to 1) for 10 examples per setting. - = the setting is not applicable for a particular level. $\bigtriangleup$ = the prompt was too large to fit the context window of the model, hence the setting was ommitted. ** = the model did not return any solution and refused to answer the question. Note that for problem 3.1, the random baseline is 0.5, since the problem solution is binary (True/False).
@@
III-A Preliminary technical limitations
III-B Overall observations from Problem 1 and Problem 2
Model specific observations
III-C Weighted vs unweighted graph traversal
III-D Effect of k-shot prompting
III-E Effect of increased graph order and jumbling node order
III-F Positive response bias within LLMs
III-G Performance of LLMs over special problems
IV PathCompare prompting
Keeping these requirements in mind, we propose a novel prompting technique specifically for graph reasoning (traversal) tasks, which we refer to as the PathCompare technique. By simply adding the line “Let’s list down all the possible paths from node {X} to node {Y}, and compare the cost to get the answer.”, we observe a significant improvement in the graph reasoning performance of all LLMs. Through this prompt, we shape an LLM response to first list down multiple paths between two specified nodes. Once multiple paths are enumerated, the task boils down to a retrieval-cum-comparison task, which increases the probability of a more accurate response.
TABLE IV: Evaluation of models on O(10) graphs with Chain-of-Thought Reasoning. The values are represented as A/B, where A depicts the average normalized binary accuracy (ranging from 0 to 1) for 10 examples per setting, and B depicts the average normalized value for partial credit for 10 examples per setting. * = a larger context variant of the corresponding LLM was used. $\bigtriangleup$ = the prompt was too large to fit the context window of the model, hence the setting was ommitted. ** = the model did not return any solution and refused to answer the question.
@@ IV PathCompare prompting
TABLE V: Evaluation of models on O(10) graphs using the PathCompare method. The values are represented as A/B, where A depicts the average normalized binary accuracy (ranging from 0 to 1) for 10 examples per setting, and B depicts the average normalized value for partial credit for 10 examples per setting. * = a larger context variant of the corresponding LLM was used. $\bigtriangleup$ = the prompt was too large to fit the context window of the model, hence the setting was omitted. ** = The model did not return any solution and refused to answer the question.

### v2

# Can LLMs perform structured graph reasoning tasks?

@@ Abstract
Pretrained Large Language Models (LLMs) have demonstrated various reasoning capabilities through language-based prompts alone, particularly in unstructured task settings (tasks purely based on language semantics). However, LLMs often struggle with structured tasks, because of the inherent incompatibility of input representation. Reducing structured tasks to uni-dimensional language semantics often renders the problem trivial. Keeping the trade-off between LLM compatibility and structure complexity in mind, we design various graph reasoning tasks as a proxy to semi-structured tasks in this paper, in order to test the ability to navigate through representations beyond plain text in various LLMs. Particularly, we design 10 distinct problems of graph traversal, each representing increasing levels of complexity, and benchmark 5 different instruct-finetuned LLMs (GPT-4, GPT-3.5, Claude-2, Llama-2 and Palm-2) on the aforementioned tasks. Further, we analyse the performance of models across various settings such as varying sizes of graphs as well as different forms of k-shot prompting. We highlight various limitations, biases and properties of LLMs through this benchmarking process, such as an inverse relation to the average degrees of freedom of traversal per node in graphs, the overall negative impact of k-shot prompting on graph reasoning tasks, and a positive response bias which prevents LLMs from identifying the absence of a valid solution. Finally, we introduce a new prompting technique specially designed for graph traversal tasks (PathCompare), which demonstrates a notable increase in the performance of LLMs in comparison to standard prompting techniques such as Chain-of-Thought (CoT).  [^1] The code for reproducing the results of this paper can be found at https://github.com/PalaashAgrawal/LLMGraphTraversal
Keywords:
Large Language Models LLM reasoning graph reasoning LLM benchmark
1 Introduction
Large language Models (LLMs) exhibit interesting reasoning abilities in various types of reasoning tasks [23], including arithmetic [27], logical [26], semantic [14], and symbolic [25]. Building on top of just raw reasoning abilities, various techniques, such as chain-of-thought (CoT) and self-consistency[22], have been proposed to improve the reasoning abilities of LLMs by enhancing the implicit knowledge in the output. However, in relative terms, graph reasoning is not well explored in the domain of LLMs, even though many important subproblems arise in graph reasoning, including subgraph matching [5] and shortest path connectivity [6]. One of the major challenges of grounding graphs through language is rooted in the fact that graphs inherently cover non-linear connectivity between multi-dimensional elements in a condensed format, and hence, reasoning over graphs requires simultaneous tracking of the state of multiple nodes. Broadly in the context of LLMs, graph reasoning is a proxy for understanding how such models establish relations between different entities in a structured setting, and is worth studying, given the impressive emergent abilities observed in recent LLMs.
The major challenge for LLMs in processing graphs directly arises from the uni-dimensionality of input representations, restricting the scope of reasoning that can be achieved using LLMs [1]. Particularly, LLMs are trained on semantic text as next-token predictors, and therefore cannot process structured data representations  [^2] In this paper, we refer to structured representations as data organized according to a specific schema or format that clearly defines the type, relationship, and arrangement of data elements. This structured format allows for efficient processing using standardized algorithms and tools. In this context, graphs are categorized as structured representations by virtue of node connections. On the other hand, LLMs are designed to process unstructured text representations, in which there is a uni-dimensional semantic relationship. By default, LLMs are not equipped to process structured representations through relationship rules. such as graphs and trees directly, without having to break down the representation into verbally semantic prompts [2]. This process often constrains the selection of tasks, and/or leads to loss of complexity. Hence, efficiently capturing the true underlying structure of graphs is particularly challenging for LLMs.
In this paper, we are particularly interested in studying fundamental graph understanding and reasoning abilities of LLMs. While various previous works have studied graph reasoning using LLMs by means of simplified graph structures as enumerated node adjacency lists and connections [2][3][4], graph tasks such as traversal are inherently reduced to trivial search retrieval problems due to flattened representations. Hence, the ability of LLMs to navigate over more complex and structured representations is still not well explored. We explore this particular question in-depth in this paper, by going one step ahead of previous works. Specifically, we prompt LLMs to reason over semi-structured graph representations, instead of plainly verbalized unstructured representations like previous works, while still maintaining the language compatibility of LLMs.  [^3] In this paper’s context, semi-structured representations involve breaking down a structured representation into uni-directional text, while still maintaining some form of element relationship. In our case, adjacency matrix representations involve a tabular relation set, which requires tracking both the column and row origin of a node. This structure introduces complexity to the reasoning task. Particularly, we are interested in exploring the ability in LLMs to handle such representational complexity.
A systematic graph-reasoning benchmarking of pretrained LLMs was performed in [17], where various interesting properties were revealed. This included preliminary reasoning capabilities in LLMs in basic graph traversal problems by parsing adjacency lists as input, and the ineffectual nature of advanced prompting techniques in the context of logical deduction. In the aforementioned work, an improvement in LLM reasoning is observed as a result of custom prompting, majorly along the lines of directing the LLM to elaborate on graph connections to facilitate node traversal. However, representation of graphs through adjacency lists does not evaluate the ability of LLMs to navigate through structured representations. Moreover, the design of graph traversal tasks in [17] is limited in terms of granularity of reasoning evaluation.
@@ 1 Introduction
In order to assess the graph reasoning abilities of LLMs in a more comprehensive way, we carry out graph evaluation through a series of increasingly complex graph problems and settings. Particularly, we construct 10 distinct graph problems requiring multi-hop reasoning and tracking. This includes tree-based graph traversals, grid-based graph-traversals as well as certain special classes of problems. We evaluate five different LLMs (GPT-3.5, GPT-4, Claude-2, Llama-2 and Palm-2). While the OpenAI family of models is thoroughly investigated in various papers, this paper studies the nature of various other contenders. For example, in this paper, we demonstrate that Anthropic’s Claude-2 is a good logical reasoning model, only sub-standard to GPT-4 among the various models publically available for access. We also choose to adopt certain design choices that highlight interesting limitations and biases in the training of various models. This includes methods like jumbling the sequence of nodes, as well as testing if the models can identify the absence of a valid solution, or in other words, testing if the models are biased to responding with a non-negative response.
@@
    We benchmark five different LLM models on graph reasoning tasks. The graph reasoning tasks are carefully designed in the form of 10 graph traversal problems in increasing orders of complexity. The design of these problems not only assesses the complexity that LLMs can handle, but also highlights interesting properties, including biases introduced due to prompting.
@@
    To study graph reasoning in depth, we also introduce many variations on top of the graph traversal tasks, including node order (graph size) variation, introduction of weighted and directed connections, and jumbling of node labels. This allows us to guarantee a wide collection of tasks not previously seen by LLMs during pretraining.
@@
    We finally propose a new prompting technique (PathCompare) which improves graph traversal performance across different LLMs. We demonstrate that this prompting technique outperforms standard prompting as well as CoT prompting in the majority (i.e. $>$50%) of the designed tasks).
@@
2 Defining Graph Reasoning Tasks
@@
2.1 Defining graph levels and complexity
@@
these graphs are defined as connected undirected graphs, where any two nodes $u,v$ are connected by atmost one edge $e\epsilon\{0,R^{+}\}$, (where $R^{+}$ represents positive real number values). In the above representation $e=0$ represents no connection, $e=1$ represents a simple unweighted connection between nodes $u,v$, and all other positive values $e~{}\epsilon~{}(R^{+}|e\neq 1$) represents a weighted connection between nodes $u,v$. By definitions, no cycles exist in a tree-based graph. However, for the sake of logical grouping of tasks in a generalizable manner, we define tree-based graphs as tree-like heirarchical structures with possible exceptions (with respect to cyclic connections). The emphasis is on navigating through these arbitrary node connections to find the shortest or least costly route. Based on this, we formulate the following problems
@@
2.2 Graph generation, prompting and evaluation
@@
In order to maintain structure of graphs without explicit flattening and enumeration of nodes/connections, we represent graph structures as adjacency matrices, rather than simpler adjacency lists. This kind of structure is not only more compact, but tests the ability of LLMs to perform multi-hop reasoning, and tracking value states more rigorously. A detailed description of our base prompting method is described in the Appendix.
@@
2.3 Selection of LLMs
In order to include a broad range of models, we select five models from four families of models (in terms of architecture and training data/method). This includes – OpenAI’s GPT-3.5 and GPT-4 [21], Anthropic’s Claude-2  [^4] It should be noted that Claude-3 was not released at the time of conducting these experiments.[20] , Google’s Palm-2 [19], and Meta’s Llama-2 [18] with 13B parameters. Since LLama-2 is a raw foundational model, we use an instruction-fine tuned version (Hermes) released by Nous Research.
3 Results and Analysis
Using the prompting and evaluation technique defined above, we evaluate the various models over the various problems. We evaluate the results of graph reasoning for problem 1, problem 2, and problem 3, by calculating averaged and normalized binary accuracies across 10 examples for each setting, and go from 0 to 1 in increments of 0.1. A model response is only marked correct if it predicted the shortest/least-cost traversal path correctly and completely. Along with binary accuracies, partial accuracies are also calculated. While we use binary accuracy to filter out models with weak reasoning abilities, we use partial accuracy for more granular comparison. These results can be visualized in Fig. 2. A detailed tabulation of these results are included in the Appendix, in tables 1 through 3.
@@ 3 Results and Analysis
3.1 Preliminary technical limitations
3.2 Overall observations from Problem 1 and Problem 2
Model specific observations
3.3 Weighted vs unweighted graph traversal
3.4 Effect of k-shot prompting
3.5 Effect of increased graph order and jumbling node order
3.6 Positive response bias within LLMs
3.7 Performance of LLMs over special problems
4 PathCompare prompting
Keeping these requirements in mind, we propose a novel prompting technique specifically for graph reasoning (traversal) tasks, which we refer to as the PathCompare technique. By simply adding the line “Let’s list down all the possible paths from node {X} to node {Y}, and compare the cost to get the answer.", we observe a significant improvement in the graph reasoning performance of all LLMs (see Figure 5 for visualization and Appendix Table 5 for detailed analysis), in comparison to standard prompting and other forms of prompting such as CoT (see Appendix Table 4). Through this prompt, we shape an LLM response to first list down multiple paths between two specified nodes. Once multiple paths are enumerated, the task boils down to a retrieval-cum-comparison task, which increases the probability of a more accurate response.
5 Summary
Appendix 0.A Appendix
0.A.1 Prompting
0.A.1.1 Base Prompt
We prompt models using adjacency matrices, labeled sequentially as A,B,C…, with 0 indicating no connections between 2 nodes, and 1 indicating otherwise. The model is asked to find the shortest path node A, to the last sequential node. An example prompt is as follows.
Given is the adjacency matrix for a unweighted undirected graph containing 10 nodes labelled A to J. The value corresponding to each row M and column N represents whether there is a connection between the two nodes, where 0 means no connection.
What is the shortest path from node A to node J? Return the sequence of nodes in response.
A B C D E F G H I J
A 0 1 0 0 0 0 0 0 0 0
B 1 0 1 0 0 0 0 0 0 0
C 0 1 0 1 0 0 0 0 0 0
D 0 0 1 0 1 0 0 0 0 0
E 0 0 0 1 0 1 0 0 0 0
F 0 0 0 0 1 0 1 0 0 0
G 0 0 0 0 0 1 0 1 0 0
H 0 0 0 0 0 0 1 0 1 0
I 0 0 0 0 0 0 0 1 0 1
J 0 0 0 0 0 0 0 0 1 0
Evaluation of the output of the models was performed manually, in order to factor in the variation of the output format by different models. In our code, input/output pairs (prompt/solution pairs) are automatically generated, which allows us to create a large scale of variations. An example of the output string generated for this corresponding problem is
A -> B -> C -> D -> E -> F -> G -> H -> I -> J
0.A.1.2 Adding k-shot prompts
Particularly for k-shot prompts (1-shot and 3-shot settings), we prepend randomly generated prompts to the task, as below.
Given is the adjacency matrix for a unweighted undirected graph containing 10 nodes labelled A to J. The value corresponding to each row M and column N represents whether there is a connection between the two nodes, where 0 means no connection.
Consider some examples
Example 1: What is the shortest path from node A to node O? Return the sequence of nodes in response.
A B C D E F G H I J K L M N O
A 0 1 0 1 0 0 0 0 0 0 0 0 1 0 0
B 1 0 1 0 0 0 0 0 0 0 0 0 0 0 0
C 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0
D 1 0 0 0 1 0 0 0 1 0 0 1 0 0 0
E 0 0 0 1 0 1 0 0 0 0 0 0 0 0 0
F 0 0 0 0 1 0 1 0 0 0 0 0 0 0 0
G 0 0 0 0 0 1 0 1 0 0 0 0 0 0 0
H 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0
I 0 0 0 1 0 0 0 0 1 0 0 0 0 0 0
J 0 0 0 0 0 0 0 1 0 1 0 0 0 0 0
K 0 0 0 0 0 0 0 0 1 0 1 0 0 0 0
L 0 0 0 1 0 0 0 0 0 0 0 0 0 0 0
M 1 0 0 0 0 0 0 0 0 0 0 0 1 0 0
N 0 0 0 0 0 0 0 0 0 0 0 1 0 1 0
O 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0
Solution: A -> M -> N -> O
Given these examples, answer the following quesiton. What is the shortest path from node A to node J? Return the sequence of nodes in response.
A B C D E F G H I J
A 0 1 0 0 0 0 0 0 0 0
B 1 0 1 0 0 1 0 0 0 0
C 0 1 0 1 0 0 0 0 0 0
D 0 0 1 0 1 0 0 0 0 0
E 0 0 0 1 0 0 0 0 0 0
F 0 1 0 0 0 0 1 0 0 0
G 0 0 0 0 0 1 0 1 0 0
H 0 0 0 0 0 0 1 0 1 0
I 0 0 0 0 0 0 0 1 0 1
J 0 0 0 0 0 0 0 0 1 0
0.A.1.3 Node Jumbling
An important variant of our graph traversal tasks involves jumbling the labels of the node order. This is primarily done, in order to avoid pretraining distribution biases in the model responses. For example, consider the adjacency matrix
A B C D E F G H I J
A 0 1 0 0 0 0 0 0 0 0
B 1 0 1 0 0 0 0 0 0 0
C 0 1 0 1 0 0 0 0 0 0
D 0 0 1 0 1 0 0 0 0 0
E 0 0 0 1 0 1 0 0 0 0
F 0 0 0 0 1 0 1 0 0 0
G 0 0 0 0 0 1 0 1 0 0
H 0 0 0 0 0 0 1 0 1 0
I 0 0 0 0 0 0 0 1 0 1
J 0 0 0 0 0 0 0 0 1 0
On jumbling, the new matrix looks something like
H C D F I E G A J B
H 0 1 0 0 0 0 0 0 0 0
C 1 0 1 0 0 0 0 0 0 0
D 0 1 0 1 0 0 0 0 0 0
F 0 0 1 0 1 0 0 0 0 0
I 0 0 0 1 0 1 0 0 0 0
E 0 0 0 0 1 0 1 0 0 0
G 0 0 0 0 0 1 0 1 0 0
A 0 0 0 0 0 0 1 0 1 0
J 0 0 0 0 0 0 0 1 0 1
B 0 0 0 0 0 0 0 0 1 0
which is essentially a renaming of node labels. Correspondingly, the modified task now involves finding the shortest node between the new "first" node and the new "last" node. Unsurprisingly, we observe a drop in the performance of models when we introduce this variation, indicating atleast some effect of pretraining data distribution bias (in the form of sequential nodel labeling A,B,C,…) in graph traversal tasks.
0.A.1.4 Weighted and Directed graph representations
While weighted connections are represented by integers $\geq 1$, directed connections are represented by removing symmetricity accross the diagonal of the matrix. A simple example to illustrate this is shown below.
Given is the adjacency matrix for a weighted directed graph containing 16 nodes labelled A to P. The value corresponding to each row M and column N represents the cost of travelling between the two nodes, where 0 means no connection.
What is the least cost path from node A to node P? Return the sequence of nodes in response.
A B C D E F G H I J K L M N O P
A 0 3 0 0 0 0 0 0 0 0 0 0 0 0 0 0
B 0 0 4 0 0 0 0 0 0 0 0 0 0 0 0 0
C 0 0 0 2 0 0 0 0 0 0 0 0 0 0 0 0
D 0 0 0 0 0 0 0 2 0 0 0 0 0 0 0 0
E 2 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0
F 0 0 0 0 1 0 0 0 0 4 0 0 0 0 0 0
G 0 0 3 0 0 4 0 0 0 0 0 0 0 0 0 0
H 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0
I 0 0 0 0 0 0 0 0 0 5 0 0 2 0 0 0
J 0 0 0 0 0 0 0 0 0 0 0 0 0 2 0 0
K 0 0 0 0 0 0 1 0 0 5 0 0 0 0 0 0
L 0 0 0 0 0 0 0 0 0 0 3 0 0 0 0 0
M 0 0 0 0 0 0 0 0 0 0 0 0 0 2 0 0
N 0 0 0 0 0 0 0 0 0 0 0 0 0 0 2 0
O 0 0 0 0 0 0 0 0 0 0 2 0 0 0 0 2
P 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
@@ A B C D E F G H I J K L M N O P
Table 1: Evaluation of models on problem category 1: tree-based graph traversals. The values are represented as A/B, where A depicts the average normalized binary accuracy (ranging from 0 to 1) for 10 examples per setting, and B depicts the average normalized value for partial credit for 10 examples per setting. * = a larger context variant of the corresponding LLM was used. $\bigtriangleup$ = the prompt was too large to fit the context window of the model, hence the setting was omitted. ** = the model did not return any solution and refused to answer the question.
@@
Table 2: Evaluation of models on problem category 2: grid-based graph traversals. The values are represented as A/B, where A depicts the average normalized binary accuracy (ranging from 0 to 1) for 10 examples per setting, and B depicts the average normalized value for partial credit for 10 examples per setting. The notations are similar to Table 1.
@@
Table 3: Evaluation of models on problem category 3: special problems. The values depict the average normalized value (ranging from 0 to 1) for 10 examples per setting. - = the setting is not applicable for a particular level. $\bigtriangleup$ = the prompt was too large to fit the context window of the model, hence the setting was ommitted. ** = the model did not return any solution and refused to answer the question. Note that for problem 3.1, the random baseline is 0.5, since the problem solution is binary (True/False).
@@
Table 4: Evaluation of models on O(10) graphs with Chain-of-Thought Reasoning. The values are represented as A/B, where A depicts the average normalized binary accuracy (ranging from 0 to 1) for 10 examples per setting, and B depicts the average normalized value for partial credit for 10 examples per setting. * = a larger context variant of the corresponding LLM was used. $\bigtriangleup$ = the prompt was too large to fit the context window of the model, hence the setting was ommitted. ** = the model did not return any solution and refused to answer the question.
@@
Table 5: Evaluation of models on O(10) graphs using the PathCompare method. The values are represented as A/B, where A depicts the average normalized binary accuracy (ranging from 0 to 1) for 10 examples per setting, and B depicts the average normalized value for partial credit for 10 examples per setting. * = a larger context variant of the corresponding LLM was used. $\bigtriangleup$ = the prompt was too large to fit the context window of the model, hence the setting was omitted. ** = The model did not return any solution and refused to answer the question.